Steam enters a turbine stage with an enthalpy of 3620 kJ/kg at 75 m/s and leaves the same stage with an enthalpy of 2800 kJ/kg at 128 m/s. Calculate the work done by steam. 


W = ____ kJ/kg


Answer in 2 decimal places


1
Expert's answer
2021-08-25T18:15:29-0400

Gives

"H_1=3620kJ/kg\\H_2=2800kJ/kg\\v_1=75m/sec\\v_2=128m/sec"

"m=\frac{1}{1000}kg"

Now

"E_{in}=E_{out}"

"KE_1+H_1=KE_2+H_2+w"

"w=(H_1-H_2)+\frac{1}{2}m(v_1^2-v_2^2)"


"w=(3620-2800)+\frac{1}{2\times1000}(75^2-128^2)"

"w=820-5.379=814.62J\/kg"

"w=0.815kJ\/kg"